OpenStudy (anonymous):

Which conditional and its converse are both true? If x = 3, then x2 = 6. If x = 2, then x2 = 4. If x2 = 4, then x = 2. If x = 1, then 2x = 2. - i think its C but i dont know

OpenStudy (anonymous):

B and D are both correct... in case of C If x2 = 4, then x = +2,-2 which is not an option

OpenStudy (anonymous):

im confused

OpenStudy (mathstudent55):

Let's look at each one. A. If x = 3, then x2 = 6. Is this true?

OpenStudy (anonymous):

you;re confused about C or B and D?

OpenStudy (mathstudent55):

What is \(3^2 =\)

OpenStudy (anonymous):

c

OpenStudy (anonymous):

im confuced a bout c

OpenStudy (mathstudent55):

C. If \(x^2 = 4\), what can x =? What numbers squared = 4?

OpenStudy (anonymous):

C says If x2 = 4, then x = 2. which is incorrect \[x=\sqrt{4}=+2 , -2\]

OpenStudy (anonymous):

\[(-2)^{2}=2^{2}=4\] remember?

OpenStudy (anonymous):

ok well idk if my are just tried but it look to me that a is correct

OpenStudy (anonymous):

A can't be right... see If x = 3, then x^2=3^2=9 while A says If x = 3, then x^2 = 6.

OpenStudy (anonymous):

then which one is correct

OpenStudy (anonymous):

nm b is correct

OpenStudy (anonymous):

thank you guys

OpenStudy (mathstudent55):

You're not done.

OpenStudy (mathstudent55):

For each conditional, you need to see if both the conditional and its converse are true. Let's start with A.: If x = 3, x^2 = 6. This is false since 3^2 = 9, so if x = 3, x^2 = 9 not 6. Here the conditinal is false, so there is no need to check the converse.

OpenStudy (mathstudent55):

B. If x = 2, then x2 = 4. This conditional is true. If x = 2, 2^2 = 4. Now we need to check the converse. If x^2 = 4, then x = 2. This converse is false. If x^2 = 4, then x = -2 or +2. For B. the conditional was true but the converse was false. So B. is not the answer.

OpenStudy (mathstudent55):

C. If x2 = 4, then x = 2. If x^2 = 4, then x = -2 or +2. This conditional is false, so there is no need to check its converse.

OpenStudy (mathstudent55):

D. If x = 1, then 2x = 2. This conditional is true. If x = 1, then 2x = 2(1) = 2, or you can multiply both sides by2 to get 2x = 2. Now let's check the converse: If 2x = 2, then x = 1. We divide both sides by 2, to get: x = 1. The converse is also true. The answer is D. because that is the only conditional that is true , and its converse is also true.

OpenStudy (anonymous):

ok i get it thank you for explaining

OpenStudy (mathstudent55):

wlcm